# Differentially Private SGD Without Clipping Bias: An Error-Feedback Approach

### v1

@@ B.1 Example of constant clipping bias
By setting the dataset size as $N=C^{\prime}+1$, and the samples are $\xi_{i}=-1,\forall~{}i=1,\dots,C^{\prime}$, and $\xi_{C^{\prime}+1}=C^{\prime}$, we can verify that $f(x)$ satisfies Assumptions 3.1-3.6 with certain constants.
@@
Appendix C Experiment Details
C.1 Adam variant of DiceSGD
@@ C.2 Efficiency analysis
In our ablation study, we observe two patterns to achieve good accuracy: (1) $C_{2}$ needs to be close to $C_{1}$. (2) $C_{1}$ needs to be small. Therefore we can write $C_{1}=C_{2}=C$ and derive an automatic version of DiceSGD: in 2, $\mathbf{v}^{t}=\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}\mathrm{clip}\left(\nabla f(\mathbf{x}^{t};\xi_{i}),C\right)+\mathrm{clip}\left(\mathbf{e}^{t},C\right)$ where $\mathrm{clip}\left(x,C\right)=\frac{C}{\|x\|}$. Setting the injected noise $\sigma_{1}=\sqrt{\frac{32T\cdot 3C^{2}\log(1/\delta)}{N^{2}\epsilon^{2}}}=\frac{\sqrt{96T\log(1/\delta)}C}{N\epsilon}$ satisfies DP. Let the learning rate $\eta^{t}$ absorbs $C$:
@@
Memory: DiceSGD requires $3$ times the memory of DPSGD-GC (which has similar time/space efficiency to the standard SGD), i.e., besides the summed clipped per-sample gradient, DiceSGD requires extra memory for both summed unclipped gradient and the feedback signal $\mathbf{e}^{t}$.
@@
ul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ B.1 Example of constant clipping bias
By setting the dataset size as $N=C^{\prime}+1$, and the samples are $\xi_{i}=-1,\forall~{}i=1,\dots,C^{\prime}$, and $\xi_{C^{\prime}+1}=C^{\prime}$, we can verify that $f(x)$ satisfies Assumptions 3.1-3.5 with certain constants.
@@
Appendix C Additional experiments
C.1 Ablation study
(a) Cifar-10
(b) Cifar-100
(a) Cifar-10
(b) Cifar-100
C.2 Training GPT-2
Ablation study for GPT-2
C.3 Adam variant of DiceSGD
@@ C.4 Efficiency analysis
In our ablation study, we observe two patterns to achieve good accuracy: (1) $C_{2}$ needs to be close to $C_{1}$. (2) $C_{1}$ needs to be small. Therefore we can write $C_{1}=C_{2}=C$ and derive an automatic version of DiceSGD: in 2, $\mathbf{v}^{t}=\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}\mathrm{clip}\left(\nabla f(\mathbf{x}^{t};\xi_{i}),C\right)+\mathrm{clip}\left(\mathbf{e}^{t},C\right)$ where $\mathrm{clip}\left(x,C\right)=\frac{C}{\|x\|}$, known as the automatic clipping in Bu et al. (2024). Setting the injected noise $\sigma_{1}=\sqrt{\frac{32T\cdot 3C^{2}\log(1/\delta)}{N^{2}\epsilon^{2}}}=\frac{\sqrt{96T\log(1/\delta)}C}{N\epsilon}$ satisfies DP. Let the learning rate $\eta^{t}$ absorbs $C$: